### v2

# Distributional Reinforcement Learning with Dual Expectile-Quantile Regression

Abstract
Distributional reinforcement learning (RL) has proven useful in multiple benchmarks as it enables approximating the full distribution of returns and makes a better use of environment samples. The commonly used quantile regression approach to distributional RL – based on asymmetric $L_{1}$ losses – provides a flexible and effective way of learning arbitrary return distributions. In practice, it is often improved by using a more efficient, hybrid asymmetric $L_{1}$-$L_{2}$ Huber loss for quantile regression. However, by doing so, distributional estimation guarantees vanish, and we empirically observe that the estimated distribution rapidly collapses to its mean. Indeed, asymmetric $L_{2}$ losses, corresponding to expectile regression, cannot be readily used for distributional temporal difference learning. Motivated by the efficiency of $L_{2}$-based learning, we propose to jointly learn expectiles and quantiles of the return distribution in a way that allows efficient learning while keeping an estimate of the full distribution of returns. We prove that our approach approximately learns the correct return distribution, and we benchmark a practical implementation on a toy example and at scale. On the Atari benchmark, our approach matches the performance of the Huber-based IQN-1 baseline after $200$M training frames but avoids distributional collapse and keeps estimates of the full distribution of returns.
**footnotetext: Both authors contributed equally to the paper.
1 Introduction
Distributional reinforcement learning (RL) (Bellemare et al., 2023) aims to maintain an estimate of the full distribution of expected return rather than only the mean. It can be used to better capture the uncertainty in the transition matrix of the environment (Bellemare et al., 2017), as well as the stochasticity of the policy being evaluated, which may enable faster and more stable training by making better use of the data samples (Mavrin et al., 2019).
Non-parametric approximations of the return distribution learned by quantile regression have proven very effective in several domains (Dabney et al., 2018a; b; Yang et al., 2019), when combined with deep RL agents such as deep Q-networks (DQN) (Mnih et al., 2013) or soft actor-critic (SAC) (Haarnoja et al., 2018). They come with the major advantage of providing guarantees for the convergence of distributional policy estimation (Dabney et al., 2018b), and in certain cases, of convergence to the optimal policy (Rowland et al., 2023), all while requiring few assumptions on the shape of the return distribution and demonstrating strong empirical performance (Dabney et al., 2018a; Yang et al., 2019). However, the best-performing quantile-based agents are often obtained by replacing the original quantile regression loss function, i.e., an asymmetric $L_{1}$ loss, by an asymmetric Huber loss, i.e., a hybrid $L_{1}$-$L_{2}$ loss. By doing so, distributional guarantees vanish. Critically, we show in Section 5.2.4 that the estimated distributions collapse to their mean in practice. In this paper, we propose a different approach, based on both quantile and expectile regression, that matches the performance of Huber-based agents while preserving distributional estimation guarantees and avoiding distributional collapse in practice.
We are not the first to note that asymmetric $L_{2}$ losses, i.e., that regress expectiles of the target distribution, tend to yield degenerate estimated distributions when training agents with temporal difference learning. Rowland et al. (2019) note that expectiles of a distribution cannot be interpreted as samples from this distribution, and therefore expectiles other than the mean cannot be directly used to compute the target values in distributional temporal difference learning. Instead, they propose to generate samples from expectiles of the distribution by adding an imputation step, that requires solving a costly root-finding problem. While theoretically justified, we found this approach to be extremely slow in practice, preventing widespread use at scale. In contrast, our dual approach only requires an additional two-layer neural network and computing a quantile loss function on top of the expectile loss function, which adds close to no computational overheads when training Atari agents on modern GPUs.
Our contributions can be summarized as follows:
  • We propose a dual expectile-quantile approach to distributional dynamic programming that provably converges to the true value distribution.
  • We release implicit expectile-quantile networks (IEQN),***Available at https://anonymous.4open.science/r/ieqn. a practical implementation of our dual approach based on implicit quantile networks (Dabney et al., 2018a).
  • We show both on a toy example and at scale on the Atari-5 benchmark that IEQN (i) avoids distributional collapse, and (ii) matches the performance of the Huber-based IQN-1 approach.
2 Background
2.1 Distributional reinforcement learning
We consider an environment modeled by a Markov decision process (MDP) $(\mathcal{S},\mathcal{A},R,T,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ are a state and action space, respectively, $R(s,a)$ denotes the stochastic reward obtained by taking action $a$ in state $s$, $T(\cdot\mid s,a)$ is the probability distribution over possible next states after taking $a$ in $s$, and $\gamma$ is a discount factor. Furthermore, we write $\pi(\cdot\mid s)$ for a (potentially stochastic) policy selecting the action depending on the current state. We consider the problem of finding a policy maximizing the average discounted return, i.e.,
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right],\vspace*{-1mm}$ (1)
where $a_{t}\sim\pi(\cdot\mid s_{t})$ and $s_{t+1}\sim T(\cdot\mid s_{t},a_{t})$. We can define the action-value random variable for policy $\pi$ as $Z^{\pi}:(s,a)\mapsto\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})$, with $s_{0}=s,a_{0}=a$. We will refer to action-value variables and their estimators as $Z$-functions in the remainder. Note that the $Q$-function, as usually defined in RL (Sutton & Barto, 2018), is given by $Q^{\pi}(s,a)=\mathbb{E}\left[Z^{\pi}(s,a)\right]$. In this work, we consider approaches that evaluate policies through distributional dynamic programming, i.e., by repeatedly applying the distributional Bellman operator $\mathcal{T}^{\pi}$ to a candidate $Z$-function:
$\mathcal{T}^{\pi}Z(s_{t},a_{t})=R(s_{t}^{\pi},a_{t}^{\pi})+\gamma Z(s_{t+1}^{\pi},a_{t+1}^{\pi}).\vspace*{-1mm}$ (2)
This operator has been shown to be a contraction in the $p$-Wasserstein distance and therefore admits a unique fixed point $Z^{\pi}$ (Bellemare et al., 2017). A major challenge of distributional RL resides in the choice of representation for the action-value distribution, as well as the exact implementation of the distributional Bellman operator.
2.2 Quantile and expectile regression
Let $Z$ be a real-valued probability distribution. The $\alpha$-quantile $q_{\alpha}$ of $Z$ is defined as a value splitting the probability mass of $Z$ in two parts of weights $\alpha$ and $1-\alpha$, respectively:
$P(z\leq q_{\alpha})=\alpha.$ (3)
Therefore, the quantile function $Q_{Z}:\alpha\mapsto q_{\alpha}$ is the inverse cumulative distribution function: $Q_{Z}=F_{Z}^{-1}$. Alternatively, quantiles are given by the minimizer of an asymmetric $L_{1}$ loss:
$q_{\alpha}=\arg\min_{q}\mathbb{E}_{z\sim Z}\left[\left(\alpha\mathds{1}_{z>q}+(1-\alpha)\mathds{1}_{z\leq q}\right)\left|z-q\right|\right].$ (4)
Expectiles and the expectile function $E_{Z}:\tau\mapsto e_{\tau}$ are defined analogously, as the $\tau$-expectile $e_{\tau}$ minimizes the asymmetric $L_{2}$ loss:
$e_{\tau}=\arg\min_{e}\mathbb{E}_{z\sim Z}\left[\left(\tau\mathds{1}_{z>e}+(1-\tau)\mathds{1}_{z\leq e}\right)\left(z-e\right)^{2}\right].\vspace*{-1mm}$ (5)
2.3 Quantiles and expectiles in distributional RL
Quantile regression has been used for distributional RL in many previous studies (Dabney et al., 2018a; b; Yang et al., 2019) where a parameterized quantile function $Q_{Z}^{\theta}(s,a,\alpha)$ is trained using a quantile temporal difference loss function derived from Eq. (4), i.e., for $N$ estimated quantiles:
$\mbox{}\mathcal{L}_{Q}\left(Q_{Z}^{\theta}(s,a,\cdot)\right)\!=\!\sum_{i=1}^{N}\sum_{j=1}^{N}l_{Q}(q_{i},z_{j})\textup{ with }l_{Q}(q_{i},z_{j})\!=\!\left(\alpha_{i}\mathds{1}_{z_{j}>q_{i}}+(1-\alpha_{i})\mathds{1}_{z_{j}\leq q_{i}}\right)\left|z_{j}-q_{i}\right|,\mbox{}\vspace*{-1mm}$ (6)
where the trainable quantile values $q_{i}=Q_{Z}^{\theta}(s,a,\alpha_{i})$ are obtained by querying the quantile function at various quantile fractions $\alpha_{i}$, which can be fixed by the designer (Dabney et al., 2018b), sampled from a distribution (Dabney et al., 2018a), or learned during training (Yang et al., 2019). In quantile-based temporal difference (QTD) learning, the target samples $z_{j}$ can be obtained by querying the estimated quantile function at the next state-action pair: $z_{j}=r+\gamma Q_{Z}^{\theta}(s^{\prime},a^{\prime},\alpha_{j})$.***We can have $a^{\prime}\sim\pi(\cdot\mid s^{\prime})$, as in actor-critic algorithms, or $a^{\prime}=\arg\max_{a}\mathcal{Q}_{Z}^{\theta}(s^{\prime},a,\alpha_{j})$ as in Q-learning. This section is agnostic to that choice but we refer to (Bellemare et al., 2023) for convergence analysis in the latter case. Indeed, because the true quantile function is the inverse CDF of the action-value distribution, Dabney et al. (2018b) and Bellemare et al. (2023) showed that, among $N$-atoms representations, quantiles at equidistant fractions minimize the $1$-Wasserstein distance with the action-value distribution and that the resulting projected Bellman operator is a contraction mapping in such a distance. Rowland et al. (2023) extended these results to prove the convergence of QTD learning under mild assumptions. We refer to these studies for a more detailed convergence analysis.
In contrast, expectile-based temporal difference (ETD) learning does not allow the same training loss as the one given by Eq. (6). We first write the generic ETD loss derived from Eq. (5):
$\mbox{}\mathcal{L}_{E}\left(E_{Z}^{\theta}(s,a,\cdot)\right)\!=\!\sum_{i=1}^{N}\sum_{j=1}^{N}l_{E}(e_{i},z_{j})\textup{ with }l_{E}(e_{i},z_{j})=\left(\tau_{i}\mathds{1}_{z_{j}>e_{i}}+(1-\tau_{i})\mathds{1}_{z_{j}\leq e_{i}}\right)\left(z_{j}-e_{i}\right)^{2},\mbox{}\vspace*{-1.5mm}$ (7)
with $e_{i}=E_{Z}^{\theta}(s,a,\tau_{i})$. Here, choosing $z_{j}=r+\gamma E_{Z}^{\theta}(s^{\prime},a^{\prime},\tau_{j})$, analogously to QTD learning and non-distributional TD learning, would cause the update to approximate a different distribution because the expectile function is in general not the inverse CDF of the return distribution, meaning expectiles cannot be considered as samples from the distribution. Rowland et al. (2019) formalized this idea using the concept of Bellman-closedness, i.e., that the projected Bellman operator yields the same statistics whether it is applied to the target distribution or to the implicit distribution given by statistics of the target distribution (i.e., in our case a uniform mixture of diracs with locations given by quantiles or expectiles). We build on this this concept to theoretically analyze our proposed agent.
3 Related Work
Distributional reinforcement learning has been shown to result in several benefits – by ascribing randomness to the value of a state-action pair, an algorithm can learn more efficiently for close states and actions (Mavrin et al., 2019), as well as capture possible stochasticity in the environment (Martin et al., 2020). Multiple families of approaches have emerged.
Estimating a parameterized distribution is a straightforward approach, and has been explored from both Bayesian (Strens, 2000; Vlassis et al., 2012) and frequentist (Jullien et al., 2023) perspectives. However, this usually requires an expensive likelihood computation, as well as making a restrictive assumption on the shape of the return distribution $Z$. For instance, assuming a normal distribution when the actual distribution is heavy-tailed can yield disappointing results.
Thus, approaches based on non-parametric estimation are also used to approximate the distribution. C51 (Bellemare et al., 2017) quantizes the domain where $Z$ has non-zero density (usually in 51 atoms, hence the name), and performs weighted classification on the atoms, by computing the cross-entropy between $Z$ and $\mathcal{T}^{\pi}Z$. While C51 increases performance over non-distributional RL, it requires the user to manually set the return bounds and is not guaranteed to minimize any $p$-Wasserstein metric with the target return distribution.
Another important non-parametric approach to the estimation of a distribution is quantile regression. Quantile regression relies on the minimization of an asymmetric $L1$ loss. Estimating quantiles allows to approximate the action-value distribution without relying on a shape assumption. QR-DQN (Dabney et al., 2018b) introduced quantile regression as a way to minimize the $1$-Wasserstein metric between $Z$ and $\mathcal{T}^{\pi}Z$. ER-DQN (Rowland et al., 2019) traded the estimation of quantiles for expectiles, at the cost of a potential distribution collapse, that they prevent via a root-finding procedure. Further, implicit quantile networks (IQN) (Dabney et al., 2018a) sample and embed quantile fractions, instead of keeping them fixed, thereby improving performance. Fully parameterized quantile functions (FQF) (Yang et al., 2019) add another network generating quantiles fractions to be estimated. We build on IQN and its expectile counterpart to propose a well-performing, non-collapsing agent.
4 Method
4.1 Dual training of quantiles and expectiles
It has been suggested that expectiles are more efficient than quantiles for approximating the underlying function (Newey & Powell, 1987; Waltrup et al., 2015), but unlike quantiles, they cannot be directly used to generate proper samples of the estimated return distribution ($z_{j}$ in Eq. (7)), which are required in distributional dynamic programming. Rowland et al. (2019) proposed an imputation strategy, i.e., a way to generate samples of a distribution that matches the current set of estimated expectiles, by solving a convex optimisation problem. In our experiments, we found that applying this imputation strategy tends to drastically increase the runtime (around 25 times slower in our setup), making experimentation with such methods close to impossible for researchers with modest computing resources. In this paper, we propose to learn a functional mapping between expectiles and quantiles and use the predicted quantiles to generate samples.
We learn a single $Z$-function using expectile regression. Therefore, we have $\forall(s,a)\in\mathcal{S}\times\mathcal{A},\tau\in[0,1],\;\hat{Z}(s,a,\tau)\mathrel{\hat{=}}E_{Z(s,a)}(\tau)$, where $Z$ is the true $Z$-function we wish to estimate. Then, we note that for non-deterministic $Z(s,a)$, the expectile function at a given state-action pair $E_{Z(s,a)}\in\mathbb{R}^{[0,1]}$ is a strictly increasing and continuous function that spans the entire convex hull of the distribution’s support (Holzmann & Klar, 2016). Meanwhile, the quantile function $Q_{Z(s,a)}\in\mathbb{R}^{[0,1]}$ spans the distribution’s support. As a consequence, every quantile is an expectile, i.e., there exists a functional mapping from quantile fractions to expectile fractions. In this work, we propose to learn such a mapper ^$\hat{\mathcal{M}}(s,a,\tau)\mathrel{\hat{=}}E^{-1}_{Z(s,a)}\circ F^{-1}_{Z(s,a)}(\tau)$ using quantile regression. We now have $\forall(s,a)\in\mathcal{S}\times\mathcal{A},\tau\in[0,1],\;\hat{Z}(s,a,\hat{\mathcal{M}}(s,a,\tau))\mathrel{\hat{=}}Q_{Z(s,a)}(\tau)$.
This way, we can simply query our estimator of quantiles at the next state-action pair to yield a sound imputation step, while the parameters of the $Z$-function are learned through expectile regression. In the following, we first prove in Theorem 4.2 that our imputation strategy based on the mapper $\hat{\mathcal{M}}$ maintains approximate Bellman-closedness, i.e., that using it to perform dynamic programming yields the correct statistics in the limit of infinite estimated quantile/expectile fractions. Then, in Section 4.3, we detail a practical implementation of dual expectile-quantile RL based on implicit quantile networks that we name IEQN.
4.2 Mapped expectiles are approximately Bellman-closed
Rowland et al. (2019) introduced the notion of Bellman-closedness of a set of statistics, i.e., that there exists a Bellman operator such that these statistics can be learned by dynamic programming. They showed that quantiles at $K$ equidistant fractions are approximately Bellman-closed, as the error between the learned quantiles and the true quantiles of the return distribution vanishes with large $K$. On the other hand, expectiles are not approximately Bellman-closed in general without an appropriate imputation strategy (i.e., a procedure to generate samples from expectiles).
With our main result, Theorem 4.2, we prove that the set of mapped expectile statistics is Bellman-closed, for a simple projected Bellman operator based on our mapper. In order to establish our main theorem, we first prove the following lemma, that ensures our dual regression estimator approximates the target distribution well in the limit of an infinite number of quantile/expectile fractions:
Lemma 4.1.
Let $\tau_{k}=\frac{2k-1}{2K}$, for $k=1,\dots,K$, and let $\Pi_{\mathcal{M}}:\mathscr{P}(\mathbb{R})\rightarrow\mathscr{P}(\mathbb{R})$ be the projected dual regression estimator defined as:
$\forall\eta\in\mathscr{P}(\mathbb{R}),\;\;\Pi_{\mathcal{M}}(\eta)=\frac{1}{K}\sum_{k=1}^{K}\delta_{E_{\eta}\left(\mathrm{floor}^{K}\left(E^{-1}_{\eta}(F^{-1}_{\eta}(\tau_{k})\right)\right)}=\frac{1}{K}\sum_{k=1}^{K}\delta_{E_{\eta}\left(\frac{2\left\lfloor K\mathcal{M}(\tau_{k})\right\rfloor+1}{2K+1}\right)}\;,\vspace*{-1mm}$
where $E_{\eta}:[0,1]\rightarrow\mathbb{R}$ is the expectile function of $\eta$, $F^{-1}_{\eta}:[0,1]\rightarrow\mathbb{R}$ is the inverse CDF – i.e., the quantile function – of $\eta$, and $\mathrm{floor}^{K}(x)=\tau_{\left\lfloor Kx\right\rfloor}$. Let $\eta,\eta_{1},\eta_{2}\in\mathscr{P}(\mathbb{R})$ be finite-support probability distributions with a finite number of discontinuities, and let $W_{1}$ be the $1$-Wasserstein distance. Then:
  1. (i)
    $\lim_{K\to\infty}W_{1}(\Pi_{\mathcal{M}}\eta,\eta)=0$, and
  2. (ii)
    $\lim_{K\to\infty}W_{1}(\Pi_{\mathcal{M}}\eta_{1},\Pi_{\mathcal{M}}\eta_{2})\leqslant W_{1}(\eta_{1},\eta_{2})$.
The proof is in Appendix C. It relies on several properties of the expectile function including its absolute continuity under our assumptions that we prove as an auxilliary result. Reusing the notation from this lemma, we can define our dual expectile-quantile projected Bellman operator. Let $\pi\in\mathscr{P}(\mathcal{A})^{\mathcal{S}}$ be a policy, we have:
$\mathcal{T}_{\mathcal{M}}^{\pi}=\Pi_{\mathcal{M}}\mathcal{T}^{\pi}\;,$ (8)
where $\mathcal{T}^{\pi}:Z(s_{t}^{\pi},a_{t}^{\pi})\mapsto R(s_{t}^{\pi},a_{t}^{\pi})+\gamma Z(s_{t+1}^{\pi},a_{t+1}^{\pi})$ is the distributional Bellman operator (see Section 2.1). We can now derive the main theorem, analogously to (Rowland et al., 2019, Theorem 4.6):
Theorem 4.2.
Consider the class of MDPs $\mathscr{M}$ with a fixed discount factor $\gamma\in[0,1)$ and immediate reward distributions with finite support and a finite number of discontinuities. Then the collection of mapped expectile (i.e., approximate quantile) statistics $s_{k}(\eta)=E_{\eta}\left(\frac{2\left\lfloor K\mathcal{M}\left(\frac{2k+1}{2K}\right)\right\rfloor+1}{2K}\right)$, for $k=1,\dots,K$ and $\eta\in\mathscr{P}(\mathbb{R})$ is approximately Bellman-closed for $\mathscr{M}$. That is, for every MDP $M=(\mathcal{S},\mathcal{A},R,T,\gamma)\in\mathscr{M}$ and policy $\pi\in\mathscr{P}(\mathcal{A})^{\mathcal{S}}$, we have
$\lim_{K\to\infty}\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\frac{1}{K}\sum_{k=1}^{K}\left|s_{k}(Z^{\pi}(s,a))-\hat{s}_{k}(s,a)\right|=0\;,\vspace*{-1mm}$
where $\hat{s}_{k}$ are the statistics learnt by repeatedly applying our operator $\mathcal{T}_{\mathcal{M}}^{\pi}$ and $Z^{\pi}$ is the true return distribution of policy $\pi$.
The proof for our main theorem can also be found in Appendix C. Theorem 4.2 guarantees that our proposed operator $\mathcal{T}_{\mathcal{M}}^{\pi}$ can be safely used in distributional RL, and we now present a practical implementation of an agent using this approach.
4.3 A practical implementation: IEQN
We use the principle described in Section 4.1 to implement IEQN (Algorithm 1), a new distributional RL agent based on implicit quantile networks (IQN) (Dabney et al., 2018a). The $Z$-function is modeled as a neural network inputting a state and a fraction $\tau\sim\mathcal{U}(0,1)$, and outputting $\tau$-expectile values for all actions. Its parameters are learned via an asymmetric $L_{2}$ loss, i.e., expectile regression. We also use a neural network to implement the mapper between quantile fractions and expectile fractions, and learn its parameters via an asymmetric $L_{1}$ loss, i.e., quantile regression.
Algorithm 1 Implicit expectile-quantile networks (IEQN) update
$Z$-function $Z_{\theta}$, mapper $m_{\phi}$, fractions $(\tau_{i})_{i=1,\dots,N}\sim\mathcal{U}([0,1])$, learning rate $\lambda$.
Collect experience $(s,a,r,s^{\prime})$
for $i=1,\dots,N$ do
     Compute expectile values $e_{i}\leftarrow Z_{\theta}(s,a,\tau_{i})$ and quantile values $q_{i}\leftarrow Z_{\theta}(s,a,m_{\phi}(\tau_{i}))$
     Compute the greedy next-action $a^{\prime}=\max_{b\in\mathcal{A}}\frac{1}{N}\sum_{i=1}^{N}Z_{\theta}(s^{\prime},b,\tau_{i})$
     Compute target samples $z_{i}\leftarrow r+\gamma\cdot\mathrm{stop\_grad}(Z_{\theta}(s^{\prime},a^{\prime},m_{\phi}(\tau_{i})))$
end for
Compute expectile loss $\mathcal{L}_{E}\leftarrow\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\left(\tau_{i}\mathds{1}_{z_{j}>e_{i}}+(1-\tau_{i})\mathds{1}_{z_{j}\leq e_{i}}\right)\left(z_{j}-e_{i}\right)^{2}$
Compute quantile loss $\mathcal{L}_{Q}\leftarrow\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\left(\tau_{i}\mathds{1}_{z_{j}>q_{i}}+(1-\tau_{i})\mathds{1}_{z_{j}\leq q_{i}}\right)\left|z_{j}-q_{i}\right|$
Update expectile function parameters $\theta\leftarrow\theta-\lambda\nabla_{\theta}\mathcal{L}_{E}$
Update mapper parameters $\phi\leftarrow\phi-\lambda\nabla_{\phi}\mathcal{L}_{Q}$
5 Experiments
We first demonstrate on a toy MDP the benefits of learning quantiles and expectiles together. We then describe our experimental setup and results on the Atari Arcade Learning Environment (ALE).
5.1 Chain MDP: A toy example
(a) Approximating a distribution with separate and dual training.
(b) Tabular distributional RL with separate and dual training.
Figure 1: (a) Approximating a bimodal distribution with quantile and expectile regression. Quantile regression approximates the inverse CDF, albeit with high variance, especially on extreme values (left, blue curves). Expectiles converge very quickly to the expectile function (left, red curves). When training a mapper to generate quantiles from expectiles, quantile estimation becomes much more efficient (right). (b) Distributional RL with function approximation in a chain MDP with 4 states, and a bimodal reward distribution at the last state. The expectile function collapses as the temporal difference error propagates to previous states (left, red curves) while the quantile function is a poor approximation of the inverse CDF (left, blue curves). Our dual method solves both problems (right).
We start by observing the effect of our proposed operator in a toy environment. The MDP comprises $4$ states, each pointing to the next through a unique action and without accumulating any reward, until the last state $s_{4}$, where the episode terminates and the agent obtains a reward sampled from a bimodal distribution $r\sim\left(\frac{1}{2}\mathcal{N}(-2,1)+\frac{1}{2}\mathcal{N}(+2,1)\right)$ (see Appendix B for a visual description).
Figure 0(a) highlights the advantageous properties of expectile regression that were introduced in prior work (Philipps, 2021a; b; Waltrup et al., 2015). When trying to approximate the distribution of terminal rewards directly from samples (left), expectile regression yields more accurate estimates than quantile regression in the low-data regime (recall that the quantile function is the inverse CDF while the expectile function is in general not). Interestingly, coupling expectile regression with our mapper (right) allows us to recover the quantile function much more efficiently than quantile regression itself.
In Figure 0(b) (left), we illustrate the deficiencies of regular quantile and expectile dynamic programming. Quantile function learning is sample-inefficient and fails to approximate the distribution within the given evaluation budget. However, we observe that the distribution information is propagated correctly through temporal difference updates, since the quantile functions estimated at each state coincide. In contrast, the expectile function collapses to the mean as the error propagates from $s_{4}$ to $s_{1}$. This is due to the fact that expectile values at the next state-action pair cannot be used as pseudo-samples of the return distribution $Z(s_{t+1})$ (Rowland et al., 2019). Finally, Figure 0(b) (right) shows that our dual training method, where the pseudo-samples of $Z(s^{\prime},a^{\prime})$ are the estimated quantiles $Z_{\theta}(s_{t+1},m_{\phi}(\tau))$, solves both issues: the expectile function does not collapse anymore and the quantile function approximation is more accurate.
5.2 Experiments on the Atari Arcade Learning Environment
5.2.1 Baselines
We experimented with the following baselines to evaluate our approach:
IQN-0, IQN-1
We approximate quantiles using the general approach described in IQN (Dabney et al., 2018a), respectively without and with a Huber loss.
IEN-Naive
We use a similar approach as for IQN, but trained with an expectile loss and a naive imputation step as described in (Rowland et al., 2019), i.e., expectile values are used as target for the temporal difference loss. The solver-based implementation described by the authors was too slow on our setup, as it was approximately 25 times slower than the other baselines.
5.2.2 Environments
We opted to conduct our experiments with the Atari Learning Environment (ALE) (Bellemare et al., 2013), following the setup of Machado et al. (2018), notably including a $25\%$ chance to perform a sticky action at each step, i.e., repeating the latest action instead of using the action predicted by the agent. This creates stochasticity in the environment, which should be captured by distributional RL agents. In order to accommodate for limited computing resources, we constrained ourselves to the Atari-5 subbenchmark (Aitchison et al., 2023), yet using 5 seeds to reduce the uncertainty in our results. We perform 25 validation episodes every $1$M steps to generate our performance curves. As is common with ALE, we report human-normalized scores, rather than raw game scores, and we aggregate them using the interquartile mean (IQM), as it is a better indicator of overall performance (compared to sample median) (Agarwal et al., 2021), due to its robustness to scale across tasks and to outliers. It is especially needed, as the presence of sticky actions increase the amount of outlier seeds.
5.2.3 Implementation details
We base all baselines and our method on the same underlying neural network, implemented in JAX (Bradbury et al., 2018). Its architecture follows the structure detailed in  Dabney et al. (2018a). We used the training loop composition of CleanRL (Huang et al., 2022). Hyperparameters can be found in Appendix A.1. We implemented the $Z$-function for all agents as a feed-forward neural network with layer normalization. We did not use the fraction proposal network introduced with FQF (Yang et al., 2019), as our method can be seen as complementary to it, and we focus on the effect of the choice of statistics. Finally, we found that using layer normalization increased performance for both our method and baselines.
As described in Algorithm 1, we only use the expectile loss to update the $Z$-function for our agent, while we use the quantile loss to update our mapper. The mapper is implemented as a two layer, residual fully-connected neural network with ReLU and Tanh activations. Since it is queried to obtain both the candidate and target values, we use a mapper-specific target network updated less frequently than the live network, using Polyak averaging (Polyak & Juditsky, 1992) with a weight of $0.5$. We share the parameters across all states, to simplify its architecture. We detail the implications of this choice in Appendix A.3.
Figure 2: Interquartile mean of the human normalized score of distributional RL agents on the Atari5 benchmark with 5 random seeds per environment. Shaded areas correspond to the 25-th and 75-th percentiles of a bootstrap distribution. A rolling average with window size of $20M$ frames is performed to enhance readability.
5.2.4 Results
In this section, we verify that our dual approach also provides benefits at scale, on a classic benchmark.
We first present in Figure 2 the aggregated results over 5 seeds on the Atari5 benchmark. We can see that despite a slower start, IEQN ends up matching the performance of IQN-1. To get statistically stronger results, we also performed a bootstrap hypothesis test on the difference of IQMs at the end of training (we average scores from the last 5 validation epochs to be robust to instabilities). We found that our method surpasses the performance of both the quantile approach (achieved significance level $0.0117$), and naive expectile approach (achieved significance level 00), thereby demonstrating the benefits of dual regression over single regression of either quantiles or expectiles.
Table 1: Average and standard deviation of the distance between quantile (respectively expectile) $0.1$ and $0.9$, relatively to the scale of the Q-function, at the end of training.
Quantiles spread Expectiles spread
IQN-0 1,25 $\pm$ 0,198 -
IQN-1 0,144 $\pm$ 0,072 -
IEN-Naive - 0,174 $\pm$ 0,195
IEQN 0,721 $\pm$ 0,142 0,465 $\pm$ 0,086
Furthermore, we verify in Table 1 that IEQN avoids distributional collapse in practice. In fact, while IQN-1’s estimated distribution is much narrower than IQN-0’s – a confirmation that the Huber loss causes distributional collapse –, IEQN’s quantile spread much larger than IQN-1’s. Moreover, the expectile spread of IEQN is much larger and stable than that of IEN-Naive, suggesting that expectile distributional RL yields degenerate distributions, as noted in (Rowland et al., 2019), but that dual expectile-quantile distributional RL avoids this collapse.
6 Conclusion
We proposed a statistics-based approach to distributional reinforcement learning that uses the simultaneous estimation of quantiles and expectiles of the action-value distribution. This approach presents the advantage of leveraging the efficiency of the expectile-based loss for both expectile and quantile estimation while solving the theoretical shortcomings of expectile-based distributional reinforcement learning, which often lead to a collapse of the expectile function in practice.
We showed on a toy environment how the dual optimization affects the statistics recovered in distributional RL: in short, the quantile function is estimated more accurately than with vanilla quantile regression and the expectile function remains consistent after several steps of temporal difference training. We then benchmarked our approach at scale, on the Atari5 benchmark. Our model, IEQN, matched the performance of the Huber-based IQN-1 and surpassed that of both expectile and quantile-based agents, demonstrating its effectiveness in practical scenarios.
We open possibilities for future research to use a distributional approach that performs well and does not collapse. For future work, we plan to investigate how the dual approach can be used in risk-aware decision-making problems, and how it performs when the goal is to optimize risk metrics such as (conditional) value-at-risk. Moreover, we plan to gather insights into what type of behavior is favored by the quantile and expectile loss, respectively.
Appendix A Hyperparameters, code and implementation details
A.1 Hyperparameters
We use JAX (Bradbury et al., 2018) to train our models. A full training procedure of $200$M training frames and corresponding validation epochs takes approximately $50$ hours in our setup.
Key Value
Discount factor $0.99$
Batch size $32$
Fraction distribution $\mathcal{U}([0,1])$
Learning rate $1\mathrm{e}^{-4}$
Random frames before training $200000$
Size of convolutional layers $[32,64,64]$
Size of fully-connected layer $512$
Critic updates per sample $2$
Buffer size $61\mathrm{e}$ 6
Frames between target network updates $35000$
Target network update rate $1.0$
Table 2: $Z$-function hyperparameters
Key Value
Layer size $64$
Learning rate $7\mathrm{e}^{-5}$
Target network update rate $0.5$
Table 3: Mapper hyperparameters
A.2 Code
Our training and evaluation loop is based on CleanRL (Huang et al., 2022). The anonymized code base is available on https://anonymous.4open.science/r/ieqn.
A.3 Sharing the mapper’s parameters
Sharing the mapper’s parameters across states and actions allows to lighten the computational burden, as it is part of the goal of this paper. We found this technique to work well in practice on the Atari5 benchmark, although it requires additional assumptions in theory. We review these assumptions in this section.
Yao & Tong (1996) showed that there exists such a shared mapping between quantile and expectile a when the regression follows a location-scale model, i.e., for random variables $X$ and $Y$:
$Y=\mu(X)+\sigma(X)\varepsilon\;,$
where $\mu$ and $\sigma$ are continuous functions, $\varepsilon$ is centered and finite-variance, and $\varepsilon,X$ are independent. When the return distribution follows this model, $X$ being the state-action variable in this context, sharing the mapper’s parameters is theoretically valid. While this may seem limiting, it does not require all state-action pairs to be allocated the same distributions, only that they share a common shape. Moreover, the location-scale family is quite broad, as it includes, e.g., Normal, Student, Cauchy, GEV distributions, and more.
In many distributional reinforcement learning scenarios, the assumption may be satisfied. For instance, when the environmental stochasticity emerges from small, independent perturbations, i.e., normally-distributed errors, the return distribution at every state will still be normally distributed as convolutions of Gaussian distributions are also Gaussian. On the other hand, this assumption can fail under high-frequency transition distributions, i.e., branching behaviors, where the same state-action pair can yield drastically different outcomes and the reward-next-state distribution has non-continuous support. We leave for future work the investigation of when sharing the mapper’s parameters across state-action pairs fails in practice.
Appendix B Toy Markov decision process
Figure 3: Toy Markov decision process
B.1 Expectile regression
Expectiles were originally introduced as a family of estimators of location parameters for a given distribution, to palliate possible heteroskedasticity of the error terms in regression (Newey & Powell, 1987; Philipps, 2021a).
Expectiles can be seen as mean estimators under missing data (Philipps, 2021b). Unlike quantiles, they span the entire convex hull of the distribution’s support, and on this ensemble, the expectile function is strictly increasing: an expectile fraction is always associated to a unique value. Expectiles have been used in reinforcement learning successfully before (Rowland et al., 2019), but in a way that requires a slow optimization step to achieve satisfactory performance. Moreover, expectile regression is subject to the same crossing issue as quantiles, albeit empirically less so (Waltrup et al., 2015). Expectiles have also been used in offline reinforcement learning to compute a soft maximum over potential outcomes seen in the offline data (Kostrikov et al., 2022).
Appendix C Proof of approximate Bellman-closedness
Our proof of Lemma 4.1 requires the absolute continuity of the expectile function. Therefore, we first prove the following lemma:
Lemma C.1.
Let $Z$ be a random variable taking values in $[a,b]$ and whose CDF admits at most a finite number of discontinuities. Then, the expectile function $E_{Z}:\tau\mapsto\arg\min_{e}\mathbb{E}_{z\sim Z}\left[\left(\tau\mathds{1}_{z>e}+(1-\tau)\mathds{1}_{z\leq e}\right)\left(z-e\right)^{2}\right]$ is absolutely continuous on $[0,1]$.
Proof.
Our proof relies on the Banach-Zarecki theorem (Zaretsky, 1925), which states that any real-valued function $f$ defined on a real bounded closed interval is absolutely continuous if and only if on this interval:
  1. (i)
    $f$ is continuous;
  2. (ii)
    $f$ has bounded variation;
  3. (iii)
    $f$ follows the Luzin N property (Luzin, 1915), i.e., the image by $f$ of a set with null Lebesgue measure also has null Lebesgue measure.
It is well-known that the expectile function is continuous on $[0,1]$ (Holzmann & Klar, 2016; Philipps, 2021b). Therefore, (i) is satisfied.
$E_{Z}$ is monotonically increasing and takes values in the finite support of $Z$. Therefore it has bounded variation and (ii) is satisfied.
In order to prove (iii), we first note that any function that is differentiable on a co-countable set has the Luzin N property (Luzin, 1915). We therefore use our assumption that $Z$ admits a finite number of discontinuities in the following.
Let $F_{Z}$ be the CDF of $Z$ and $D=\left\{z\in[a,b]\;,\;\lim_{x\to z}F_{Z}(x)\neq F_{Z}(z)\right\}$ be the finite set of points at which $F_{Z}$ is not continuous. $D$ is a finite set within a metric space and therefore closed. As a consequence, its complement $C_{[a,b]}=[a,b]\backslash D$ is open in $[a,b]$, i.e., $\forall z\in C_{[a,b]},\exists\varepsilon>0$ such that $\forall x\in[a,b]\;,\;d(x,z)<\varepsilon\Rightarrow x\in C_{[a,b]}$. In other words, if $F_{Z}$ is continuous at a point within $[a,b]$, it is also continuous in a neighborhood of that point within $[a,b]$. By assumption, the set $C^{\mathcal{N}}_{[a,b]}=\left\{z\in[a,b]\;,\;\exists\varepsilon>0\;,\;\forall x\in[a,b]\;,\;d(x,z)<\varepsilon\Rightarrow x\in C_{[a,b]}\right\}$ of points where $F_{Z}$ is continuous in a neighborhood of said point is therefore co-finite.
It has been shown that the expectile function $E_{Z}$ is continuously differentiable at any point $\tau\in[0,1]$ such that $F_{Z}$ is continuous in a neighborhood of $E_{Z}(\tau)$ (Holzmann & Klar, 2016; Newey & Powell, 1987). The expectile function is bijective (Philipps, 2021b) so the set of points where $E_{Z}$ is differentiable $\mathcal{D}_{[a,b]}^{E_{Z}}=E_{Z}^{-1}\left(C^{\mathcal{N}}_{[a,b]}\right)$ is also a co-finite set.
The expectile function is differentiable on a co-finite (and thus co-countable) set, i.e., it has the Luzin N property (Luzin, 1915), which yields (iii).
We can finally apply the Banach-Zarecki theorem and conclude that the expectile function $E_{Z}$ is absolutely continuous on $[0,1]$. ∎
We can now use the absolute continuity of the expectile function under our assumptions to prove the following lemma:
See 4.1
Proof.
We first prove (i). Thanks to the triangle inequality, we have :
$\begin{split}W_{1}(\Pi_{\mathcal{M}}\eta,\eta)\leqslant W_{1}(\Pi_{\mathcal{M}}\eta,\Pi_{Q}\eta)+W_{1}(\Pi_{Q}\eta,\eta)\;,\end{split}$ (9)
where $\Pi_{Q}$ is the projected quantile regression estimator defined as:
$\forall\eta\in\mathscr{P}(\mathbb{R}),\;\;\Pi_{Q}(\eta)=\frac{1}{K}\sum_{k=1}^{K}\delta_{F^{-1}_{\eta}(\tau_{k})}\;.$
Rowland et al. (2019, Lemma 3.2) showed that we have $W_{1}(\Pi_{Q}\eta,\eta)=\mathcal{O}\left(\frac{1}{K}\right)$. We now turn to the first term:
$\begin{split}W_{1}(\Pi_{\mathcal{M}}\eta,\Pi_{Q}\eta)&=\sum_{i=0}^{K-1}\frac{1}{K}\left|E_{\eta}\left(\mathrm{floor}^{K}\left(E^{-1}_{\eta}\left(F^{-1}_{\eta}\left(\frac{2i+1}{2K}\right)\right)\right)\right)-F^{-1}_{\eta}\left(\frac{2i+1}{2K}\right)\right|\\ &=\sum_{i=0}^{K-1}\frac{1}{K}\left|E_{\eta}\left(\mathrm{floor}^{K}\left(E^{-1}_{\eta}\left(F^{-1}_{\eta}\left(\frac{2i+1}{2K}\right)\right)\right)\right)-E_{\eta}\left(E^{-1}_{\eta}\left(F^{-1}_{\eta}\left(\frac{2i+1}{2K}\right)\right)\right)\right|\\ &\leqslant\sum_{i=0}^{K-1}\frac{1}{K}\left|E_{\eta}\left(\mathrm{floor}^{K}\left(E^{-1}_{\eta}\left(F^{-1}_{\eta}\left(\frac{2i+1}{2K}\right)\right)\right)\right)-{}\right.\\ &\hskip 142.26378pt\left.E_{\eta}\left(\mathrm{floor}^{K}\left(E^{-1}_{\eta}\left(F^{-1}_{\eta}\left(\frac{2i+1}{2K}\right)\right)\right)+\frac{1}{K}\right)\right|\;,\end{split}$ (10)
where the last inequality is obtained thanks to the monotonicity of the expectile function. By absolute continuity of the expectile function under our assumptions (proven in Lemma C.1), we have:
$\lim_{K\to\infty}\left|E_{\eta}\left(\mathrm{floor}^{K}\left(E^{-1}_{\eta}\left(F^{-1}_{\eta}\left(\frac{2i+1}{2K}\right)\right)\right)\right)-E_{\eta}\left(\mathrm{floor}^{K}\left(E^{-1}_{\eta}\left(F^{-1}_{\eta}\left(\frac{2i+1}{2K}\right)\right)\right)+\frac{1}{K}\right)\right|=0,$ (11)
from which we can deduce $\lim_{K\to\infty}W_{1}(\Pi_{\mathcal{M}}\eta,\Pi_{Q}\eta)=0$ and finally $\lim_{K\to\infty}W_{1}(\Pi_{\mathcal{M}}\eta,\eta)=0$.
(ii) results from (i) and the triangle inequality:
$W_{1}(\Pi_{\mathcal{M}}\eta_{1},\Pi_{\mathcal{M}}\eta_{2})\leqslant W_{1}(\Pi_{\mathcal{M}}\eta_{1},\eta_{1})+W_{1}(\eta_{1},\eta_{2})+W_{1}(\eta_{2},\Pi_{\mathcal{M}}\eta_{2})$ (12)
so
$\lim_{K\to\infty}W_{1}(\Pi_{\mathcal{M}}\eta_{1},\Pi_{\mathcal{M}}\eta_{2})\leqslant W_{1}(\eta_{1},\eta_{2}).\qed\mbox{}$
We now use this lemma in the proof of the main theorem: See 4.2
Proof.
We define $Z$, the fixed point of our operator $\mathcal{T}^{\pi}_{\mathcal{M}}=\Pi_{\mathcal{M}}\mathcal{T}^{\pi}$. Thanks to Lemma 4.1, we have:
$\begin{split}\lim_{K\to\infty}\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\frac{1}{K}\sum_{k=1}^{K}\left|s_{k}(Z^{\pi}(s,a))-\hat{s}_{k}(s,a)\right|&=\lim_{K\to\infty}\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}W_{1}(\Pi_{\mathcal{M}}Z(s,a),\Pi_{\mathcal{M}}Z^{\pi}(s,a))\\ &\overset{4.1(i)}{\leqslant}\lim_{K\to\infty}\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}W_{1}(Z(s,a),Z^{\pi}(s,a))\;.\end{split}$ (13)
Then,
[…]
$\hskip 113.81102ptW_{1}(\Pi_{\mathcal{M}}Z^{\pi}(s,a),Z^{\pi}(s,a))$ (15)
=$\overset{4.1(i)}{=}$ $\lim_{K\to\infty}\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}W_{1}(Z(s,a),\Pi_{\mathcal{M}}Z^{\pi}(s,a))$ (16)
[…]
$\overset{4.1(ii)}{\leqslant}$ $\lim_{K\to\infty}\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}W_{1}(\mathcal{T}^{\pi}Z(s,a),\mathcal{T}^{\pi}Z^{\pi}(s,a))$ (18)
[…]
where the last line follows from (Bellemare et al., 2017, Lemma 3), i.e., that the distributional Bellman operator is a $\gamma$-contraction in Wasserstein distance.
Finally, we have:
$\lim_{K\to\infty}\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\frac{1}{K}\sum_{k=1}^{K}\left|s_{k}(Z^{\pi}(s,a))-\hat{s}_{k}(s,a)\right|=0.\qed$